OpenStudy (lgbasallote):

the froghopper philaenus spumarius is supposedly the best jumpre in the animal kingdom. to start a jump, this insect can accelerate at 4.00 km/s^2 over a distance of 2.00 mm as it straightens its specially adapted "jumping legs". Assume the acceleration is constant. Find the upward velocity with which the insect takes off

OpenStudy (anonymous):

Use the formula \[V^2-V_{0}^2=2 a x\] a=4000 m/s^2 x=2 mm=0.002 m

OpenStudy (anonymous):

at t=0 insect is at rest so V0=0

OpenStudy (anonymous):

V is the take off velocity

OpenStudy (lgbasallote):

soyou're taking the starting point as origin..mmhmm

OpenStudy (lgbasallote):

so it'sjust v^2 = 2ax?

OpenStudy (anonymous):

yes thats right

OpenStudy (lgbasallote):

so v^2 = -2(4000)(0.002)?

OpenStudy (anonymous):

why u put the negative sign?

OpenStudy (lgbasallote):

lol woops nevermind that haha

OpenStudy (lgbasallote):

so is that right?

OpenStudy (anonymous):

yes i got v=4 m/s

OpenStudy (lgbasallote):

4 m/s?

OpenStudy (lgbasallote):

shouldnt it be 8

OpenStudy (anonymous):

v^2=16 and v=4 ...

OpenStudy (anonymous):

yup, i also got 4.

OpenStudy (lgbasallote):

oops forgot to multiply by 2...thanks

OpenStudy (lgbasallote):

then take the square root